# Equivariant Ehrhart theory, commutative algebra and invariant triangulations of polytopes

### v2

@@ 3.3. Translative actions
Let $G=\mathbb{Z}/5\mathbb{Z}=\langle\sigma\rangle$, and let $M=\mathbb{Z}^{5}/\mathbb{Z}(e_{1}+\cdots+e_{5})$. Let $\bar{e}_{i}$ be the image of $e_{i}$ in $M$ for $1\leq i\leq 5$. Assume that $G$ acts on $M$ by cyclically shifting coordinates i.e. $\sigma\cdot\bar{e}_{i}=\bar{e}_{i+1}$ for $1\leq i\leq 5$, where $\bar{e}_{6}=\bar{e}_{1}$ Let $\bar{f}_{i}:=\sigma^{i-1}\cdot(\bar{e}_{1}+\bar{e}_{2})$ for $1\leq i\leq 5$. Let $P$ be the convex hull of $\{\bar{e}_{i},\bar{f}_{i}:1\leq i\leq 5\}$. Then $P$ is a $4$-dimensional reflexive polytope. It has a unique interior lattice point, and its other lattice points are $\{\bar{e}_{i},\bar{f}_{i}:1\leq i\leq 5\}$, which are also the vertices of $P$. Let $Q=\operatorname{Conv}\left\{{\bar{e}_{1},\bar{e}_{3},\bar{f}_{1},\bar{f}_{2}}\right\}$. Then $Q$ is $2$-dimensional and isomorphic to $[0,1]^{2}\subset\mathbb{R}^{2}$ as a lattice polytope. The boundary of $P$ has $15$ facets in $3$ $G$-orbits. Representatives of the orbits are $F_{1}=\operatorname{Conv}\left\{{Q,\bar{f}_{3}}\right\}$, $F_{2}=\operatorname{Conv}\left\{{Q,\bar{f}_{5}}\right\}$ and $F_{3}=\operatorname{Conv}\left\{{\bar{e_{1}},\bar{e_{3}},\bar{f_{3}},\bar{f_{5}}}\right\}$. Here $F_{1}$ and $F_{2}$ are pyramids over $Q$, and $F_{3}$ is a simplex. Any $G$-invariant regular lattice polyhedral subdivision of $P$ restricts to the trivial subdivision of $Q$ and hence is not a triangulation. On the other hand, if we fix a choice of diagonal $e$ of $Q$ e.g. $e=[\bar{e}_{1},\bar{f}_{2}]$, then there is a unique $G$-invariant lattice triangulation $\mathcal{S}^{\prime}$ of the boundary of $P$ such that the restriction $\mathcal{S}^{\prime}|_{F_{i}}$ is the pulling refinement of $F_{i}$ by the vertices of $e$ for $i=1,2$, and $\mathcal{S}^{\prime}|_{F_{3}}$ is the trivial subdivision. That is, with the choice of $e$ above, the facets of $\mathcal{S}^{\prime}$ have $5$ orbits with representatives $F_{1}^{\prime}=\operatorname{Conv}\left\{{e,\bar{e}_{3},\bar{f}_{3}}\right\}$, $F_{1}^{\prime\prime}=\operatorname{Conv}\left\{{e,\bar{f}_{1},\bar{f}_{3}}\right\}$, $F_{2}^{\prime}=\operatorname{Conv}\left\{{e,\bar{e}_{3},\bar{f}_{5}}\right\}$, $F_{2}^{\prime\prime}=\operatorname{Conv}\left\{{e,\bar{f}_{1},\bar{f}_{5}}\right\}$ and $F_{3}$. We then obtain a $G$-invariant lattice triangulation $\mathcal{S}$ of $P$ by taking the join of $\mathcal{S}^{\prime}$ and the origin i.e. the facets of $\mathcal{S}$ are the convex hulls of the facets of $\mathcal{S}^{\prime}$ and the origin. In this example, $h^{*}(P,\rho;t)=1+t+t^{2}+t^{3}+t^{4}+t(1+2t+t^{2})\chi_{\operatorname{reg}}$, where $\chi_{\operatorname{reg}}$ is the regular representation of $G$.
@@
We continue with the setup of Example 2.26 and Example 3.7. That is, $\psi:\mathbb{Z}^{d+1}\to\mathbb{Z}$, $\psi(u_{1},\ldots,u_{d+1})=\sum_{i=1}^{d+1}u_{i}$, and $M^{\prime}=\psi^{-1}(\binom{d+2}{2})$ is an affine lattice in the real affine space $W=\psi_{\mathbb{R}}^{-1}(\binom{d+2}{2})$. The symmetric group $G=\operatorname{Sym}_{d+1}$ acts on the $(d+1)$-permutahedron $P\subset W$. Assume furthermore that $d+1=p$ is prime and consider the subgroup $H=\langle(1,2,\ldots,p)\rangle\cong\mathbb{Z}/p\mathbb{Z}$ of $G$. That is, $H$ acts on $P$ by cyclically shifting coordinates. When $p=2$, $d=1$ and $P$ is a simplex. Assume that $p$ is an odd prime. We claim that there exists an $H$-invariant translative regular lattice triangulation of $P$. Combined with Theorem 1.4, this gives a generalization of [EKS23, Theorem 3.52] that states that condition (2) in Theorem 1.4 holds in this case. To prove the claim, recall from Example 3.7 that the facets of $P$ are $F_{I}$ where $I=(I_{1},I_{2})$ is an ordered partition of $\{1,\ldots,p\}$, and the vertices of $F_{I}$ are precisely all $v=(v_{1},\ldots,v_{d+1})$ such that $\{v_{i}:i\in I_{1}\}=\{1,2,\ldots,|I_{1}|\}$. Then $h\cdot F_{I}\cap F_{I}=F_{h\cdot I}\cap F_{I}=\emptyset$ for all $h\neq\operatorname{id}$ in $H$, and it follows that the action of $H$ on $K=\partial P$ is translative. Let $Q$ be the interior $H$-fixed lattice point $(\frac{p+1}{2},\ldots,\frac{p+1}{2})$. The claim now follows from Corollary 3.17 and Corollary 3.18.
@@ 3.4. The two dimensional case
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. Assume that $d=2$. Then the following conditions are equivalent:
@@
The implication (1) $\Rightarrow$ (2) is clear. The implication (2) $\Rightarrow$ (3) follows from Theorem 1.4. We now verify the implication (3) $\Rightarrow$ (4). Suppose that (4) doesn’t hold. We need to show that $h^{*}(P,\rho;t)$ is not a polynomial. Firstly, consider case 4(a) above. Then the $g$-fixed locus $P^{g}$ of $P$ corresponds to the point $\{(\frac{1}{2},\frac{1}{2})\}$, and hence $\operatorname{Ehr}(P,\rho;t)(g)=\frac{1}{1-t^{2}}$. Since $\det(I-\widetilde{M}_{\mathbb{C}}t)(g)=(1-t)(1+t^{2})$, we have that $h^{*}(P,\rho;t)(g)=\frac{1+t^{2}}{1+t}$ has a pole at $t=-1$. Secondly, in case 4(b), [Sta11a, Lemma 7.3] implies that $h^{*}(P,\rho;t)$ is not a polynomial.
@@
The subgroups of $\operatorname{GL}(2,\mathbb{Z})$, up to conjugacy, are classified [Vos65]. They are precisely the subgroups of the dihedral group $D_{4}$ of symmetries of $[-1,1]^{2}$, and the subgroups of the dihedral group $D_{6}$ of symmetries of the hexagon with vertices {±$\{\pm e_{1},\pm(e_{1}+e_{2}),\pm e_{2}\}$. Consider an element $g\neq\operatorname{id}$ in $G$. By Remark 2.9, the degree $2$ characteristic polynomial of the action of $g$ on $M_{\mathbb{C}}$ is a product of cyclotomic polynomials. Hence $\operatorname{ord}(g)\in\{2,3,4,6\}$ and if $1$ is an eigenvalue of $g$ acting on $M_{\mathbb{C}}$ then $\operatorname{ord}(g)=2$ (this also follows from the classification of subgroups of $\operatorname{GL}(2,\mathbb{Z})$). If $1$ is not an eigenvalue for the action of $g$ on $M_{\mathbb{C}}$, then by Lemma 2.5 and Lemma 2.6 either $\operatorname{ord}(g)\in\{2,4\}$ and $c\in\frac{1}{2}M$ or $\operatorname{ord}(g)=3$ and $c\in\frac{1}{3}M$. If $\operatorname{ord}(g)=4$ and $c\in\frac{1}{2}M$, then we see from the classification that we are in case 4(a), a contradiction. In particular, we have shown that $\operatorname{ord}(g)\in\{2,3\}$.
@@
Let $J$ be the set of lattice points in $P$ that are not vertices of $\mathcal{S}$. Since $\mathcal{S}$ equals the pulling refinement of $\mathcal{S}$ by $J$, we conclude that $J$ is the emptyset. Hence each facet $F$ of $\mathcal{S}$ is an empty polytope, in the sense that $F\cap M$ is precisely the set of vertices of $F$. On the other hand, it is well-known and straightforward to check that the only empty $2$-dimensional lattice polytopes in $\mathbb{R}^{2}$, up to unimodular equivalence, are the standard simplex and $[0,1]^{2}$. Fix a facet $F$ of $\mathcal{S}$ that is not a simplex. Then we may assume that $M=\mathbb{Z}^{2}$ and $F=[0,1]^{2}$. Label the vertices of $F$: $u=0$, $v=e_{1}$, $w=e_{1}+e_{2}$, $x=e_{2}$.
@@ 4.1. Commutative algebra
First recall that any surjection of graded $\mathbb{C}G$-modules splits. Indeed, if $f:V\to W$ is a surjection of graded $\mathbb{C}G$-modules and $h^{\prime}:W\to V$ is a splitting of $f$ as a map of graded $\mathbb{C}$-vector spaces, then one may define $h:W\to V$ by $h(w)=\frac{1}{|G|}\sum_{g\in G}gh^{\prime}(g^{-1}w)$ for all $w$ in $W$. Consider the graded subring $S=\mathbb{C}[x_{1},\ldots,x_{n}]$ of $R$ and let $\mathfrak{m}=(x_{1},\ldots,x_{n})$. Note that each $x_{i}$ has degree $r$. By Theorem 4.4, $S$ is isomorphic to a polynomial ring in $n$ variables. Since $R$ is Cohen-Macaulay, $R$ is a free $S$-module by Theorem 4.6. Consider the surjection of graded $\mathbb{C}G$-modules $p:R\to R/J=R/\mathfrak{m}R$, and choose a splitting of graded $\mathbb{C}G$-modules $\sigma:R/J\to R$. Let $y_{1},\ldots,y_{s}$ be a $\mathbb{C}$-basis of $R/J$. Then $\sigma(y_{1}),\ldots,\sigma(y_{s})$ generate $R$ as an $S$-module by the graded version of Nakayama’s lemma (see [Eis95, Exercise 4.6a] and [AM69, Proposition 2.8]). Here $s$ is the rank of $R$ as a free $S$-module and hence $\sigma(y_{1}),\ldots,\sigma(y_{s})$ is a a basis of $R$ as an $S$-module [AM69, Exercise 3.15].
@@ 4.2. The principal $A$-determinant and nondegenerate polynomials
[GKZ94, Proposition 8.2.1] Let $A\subset M$ be a finite set with convex hull $P\subset M_{\mathbb{R}}$. With the notation above, consider the polynomial ring $\mathbb{Z}[z_{i,a}]$ for some formal variables $\{z_{i,a}:1\leq i\leq d+1,a\in A\}$. Then there is an irreducible polynomial (unique up to sign) $R_{A}(z_{i,a})\in\mathbb{Z}[z_{i,a}]$ called the $A$-resultant satisfying the following property: Let $F_{1},\ldots,F_{d+1}$ be elements of $S_{P}$ with support contained in $A\times\{1\}$. For $11\leq i\leq d+$ 1, write $F_{i}=\sum_{a\in A}\lambda_{i,a}x^{(a,1)}$ for some $\lambda_{i,a}\in\mathbb{C}$. Then $F_{1},\ldots,F_{d+1}$ is a l.s.o.p. for $S_{P}$ if and only if $R_{A}(\lambda_{i,a})\neq 0$.
@@
In Definition 4.9, the statement given is different but equivalent to that in [GKZ94, Proposition 8.2.1]. Namely, [GKZ94] associate to $A$ the semigroup algebra $S_{A}:=\mathbb{C}[C_{P}\cap\widetilde{M}_{A}]$ and the (not necessarily normal) toric variety $\operatorname{Proj}S_{A}$. With the notation of Definition 4.9, the $A$-resultant in [GKZ94] is determined by the condition that $R_{A}(\lambda_{i,a})\neq 0$ if and only if the subvariety $\operatorname{Proj}(S_{A}/(F_{1},\ldots,F_{d+1}))$ of $\operatorname{Proj}S_{A}$ is empty. We explain why this is equivalent to Definition 4.9 cf. the proof of [Bat93, Theorem 4.8] (stated in Theorem 4.18 below). First observe that $S_{A}\subset S_{P}$ and $S_{P}$ is integral over $S_{A}$. This follows, for example, by choosing a regular triangulation of $P$ with vertices in $A$, and then noting that every element in $C_{P}\cap\widetilde{M}$ is a sum of elements of $A\times\{1\}$ with nonnegative rational coefficients. In particular, $S_{A}$ has Krull dimension $d+1$ and $S$ is a finitely generated $S_{A}$-module by Lemma 4.1. Now $\operatorname{Proj}(S_{A}/(F_{1},\ldots,F_{{d+1}}))$ is empty if and only if $\operatorname{Spec}(S_{A}/(F_{1},\ldots,F_{{d+1}}))$ is 00-dimensional. The latter holds if and only if $S_{A}/(F_{1},\ldots,F_{{d+1}})$ has Krull dimension 00 i.e. $F_{1},\ldots,F_{{d+1}}$ is a l.s.o.p. for $S_{A}$. By Corollary 4.5, $F_{1},\ldots,F_{{d+1}}$ is a l.s.o.p. for $S_{A}$ if and only if $F_{1},\ldots,F_{{d+1}}$ is a l.s.o.p. for $S_{P}$.
@@
Let $\widetilde{N}_{A}=\operatorname{Hom}(\widetilde{M}_{A},\mathbb{Z})$ and $\widetilde{N}=\operatorname{Hom}(\widetilde{M},\mathbb{Z})$ be the dual lattices to $\widetilde{M}_{A}$ and $\widetilde{M}$ respectively. Then $(\widetilde{N}_{A})_{\mathbb{C}}=(\widetilde{N})_{\mathbb{C}}$ and we have the natural pairing $\langle\cdot,\cdot\rangle:\widetilde{M}_{\mathbb{C}}\times\widetilde{N}_{\mathbb{C}}\to\mathbb{C}$. Given $v$ in $\widetilde{N}_{\mathbb{C}}$, we define a $\widetilde{M}$-graded $\mathbb{C}$-linear map:
@@
[GKZ94, 10.1.A] Let $A\subset M$ be a finite set with convex hull $P\subset M_{\mathbb{R}}$. With the notation above, consider the polynomial ring $\mathbb{Z}[z_{a}]$ for some formal variables $\{z_{a}:a\in A\}$. Fix a choice of basis $v_{1},\ldots,v_{d+1}$ of $\widetilde{N}_{A}$. For $a$ in $A$ and $11\leq i\leq d+$ 1, let $w_{i,a}:=\langle(a,1),v_{i}\rangle z_{a}$. Then $E_{A}(z_{a}):=R_{A}(w_{i,a})\in\mathbb{Z}[z_{a}]$.
@@
Definition 4.12 could alternatively have been described as follows. Fix a choice of basis $v_{1},\ldots,v_{d+1}$ of $\widetilde{N}_{A}$. The principal $A$-determinant $E_{A}(z_{a})\in\mathbb{Z}[z_{a}]$ is the (unique up to sign) polynomial satisfying the following property: Let $F$ be an element of $S_{P}$ with support contained in $A\times\{1\}$, and write $F=\sum_{a\in A}\lambda_{a}x^{(a,1)}$ for some $\lambda_{a}\in\mathbb{C}$. For all $11\leq i\leq d+$ 1, write $\partial_{v_{i}}F=\sum_{a\in A}\lambda_{i,a}x^{(a,1)}$ for some $\lambda_{i,a}\in\mathbb{C}$. Then $E_{A}(\lambda_{a})=R_{A}(\lambda_{a,i})$. In particular, with this notation, $\partial_{v_{1}}F,\ldots,\partial_{v_{d+1}}F$ is a l.s.o.p. for $S_{P}$ if and only if $E_{A}(\lambda_{a})\neq 0$.
@@
where $\operatorname{Vol}(P)$ is the normalized volume of $P$ with respect to $M$ and $[\widetilde{M}:\widetilde{M}_{A}]$ is the index of the sublattice $\widetilde{M}_{A}$ of $\widetilde{M}$. Importantly, $E_{A,M}(\lambda_{a})=0$ if and only if $E_{A}(\lambda_{a})=0$.
@@
See [Stu94, (28)] for an analogous result to Theorem 4.19 for the $A$-resultant.

### v3

@@ 3.3. Translative actions
Let $G=\mathbb{Z}/5\mathbb{Z}=\langle\sigma\rangle$, and let $M=\mathbb{Z}^{5}/\mathbb{Z}(e_{1}+\cdots+e_{5})$. Let $\bar{e}_{i}$ be the image of $e_{i}$ in $M$ for $1\leq i\leq 5$. Assume that $G$ acts on $M$ by cyclically shifting coordinates i.e. $\sigma\cdot\bar{e}_{i}=\bar{e}_{i+1}$ for $1\leq i\leq 5$, where $\bar{e}_{6}=\bar{e}_{1}$ . Let $\bar{f}_{i}:=\sigma^{i-1}\cdot(\bar{e}_{1}+\bar{e}_{2})$ for $1\leq i\leq 5$. Let $P$ be the convex hull of $\{\bar{e}_{i},\bar{f}_{i}:1\leq i\leq 5\}$. Then $P$ is a $4$-dimensional reflexive polytope. It has a unique interior lattice point, and its other lattice points are $\{\bar{e}_{i},\bar{f}_{i}:1\leq i\leq 5\}$, which are also the vertices of $P$. Let $Q=\operatorname{Conv}\left\{{\bar{e}_{1},\bar{e}_{3},\bar{f}_{1},\bar{f}_{2}}\right\}$. Then $Q$ is $2$-dimensional and isomorphic to $0,1]^{2}\subset\mathbb{R}^{2}$ as a lattice polytope. The boundary of $P$ has $15$ facets in 33 $G$-orbits. Representatives of the orbits are $F_{1}=\operatorname{Conv}\left\{{Q,\bar{f}_{3}}\right\}$, $F_{2}=\operatorname{Conv}\left\{{Q,\bar{f}_{5}}\right\}$ and $F_{3}=\operatorname{Conv}\left\{{\bar{e_{1}},\bar{e_{3}},\bar{f_{3}},\bar{f_{5}}}\right\}$. Here $F_{1}$ and $F_{2}$ are pyramids over $Q$, and $F_{3}$ is a simplex. Any $G$-invariant regular lattice polyhedral subdivision of $P$ restricts to the trivial subdivision of $Q$ and hence is not a triangulation. On the other hand, if we fix a choice of diagonal $e$ of $Q$ e.g. $e=[\bar{e}_{1},\bar{f}_{2}]$, then there is a unique $G$-invariant lattice triangulation $\mathcal{S}^{\prime}$ of the boundary of $P$ such that the restriction $\mathcal{S}^{\prime}|_{F_{i}}$ is the pulling refinement of $F_{i}$ by the vertices of $e$ for $i=1,2$, and $\mathcal{S}^{\prime}|_{F_{3}}$ is the trivial subdivision. That is, with the choice of $e$ above, the facets of $\mathcal{S}^{\prime}$ have $5$ orbits with representatives $F_{1}^{\prime}=\operatorname{Conv}\left\{{e,\bar{e}_{3},\bar{f}_{3}}\right\}$, $F_{1}^{\prime\prime}=\operatorname{Conv}\left\{{e,\bar{f}_{1},\bar{f}_{3}}\right\}$, $F_{2}^{\prime}=\operatorname{Conv}\left\{{e,\bar{e}_{3},\bar{f}_{5}}\right\}$, $F_{2}^{\prime\prime}=\operatorname{Conv}\left\{{e,\bar{f}_{1},\bar{f}_{5}}\right\}$ and $F_{3}$. We then obtain a $G$-invariant lattice triangulation $\mathcal{S}$ of $P$ by taking the join of $\mathcal{S}^{\prime}$ and the origin i.e. the facets of $\mathcal{S}$ are the convex hulls of the facets of $\mathcal{S}^{\prime}$ and the origin. In this example, $h^{*}(P,\rho;t)=1+t+t^{2}+t^{3}+t^{4}+t(1+2t+t^{2})\chi_{\operatorname{reg}}$, where $\chi_{\operatorname{reg}}$ is the regular representation of $G$.
@@
We continue with the setup of Example 2.26 and Example 3.7. That is, $\psi:\mathbb{Z}^{d+1}\to\mathbb{Z}$, $\psi(u_{1},\ldots,u_{d+1})=\sum_{i=1}^{d+1}u_{i}$, and $M^{\prime}=\psi^{-1}(\binom{d+2}{2})$ is an affine lattice in the real affine space $W=\psi_{\mathbb{R}}^{-1}(\binom{d+2}{2})$. The symmetric group $G=\operatorname{Sym}_{d+1}$ acts on the $(d+1)$-permutahedron $P\subset W$.
Assume furthermore that $d+1=p$ is prime and consider the subgroup $H=\langle(1,2,\ldots,p)\rangle\cong\mathbb{Z}/p\mathbb{Z}$ of $G$. That is, $H$ acts on $P$ by cyclically shifting coordinates. When $p=2$, $d=1$ and $P$ is a simplex. Assume that $p$ is an odd prime. We claim that there exists an $H$-invariant translative regular lattice triangulation of $P$. Combined with Theorem 1.4, this gives a generalization of [EKS23, Theorem 3.52] that states that condition (2) in Theorem 1.4 holds in this case.
To prove the claim, recall from Example 3.7 that the facets of $P$ are $F_{I}$ where $I=(I_{1},I_{2})$ is an ordered partition of $\{1,\ldots,p\}$, and the vertices of $F_{I}$ are precisely all $v=(v_{1},\ldots,v_{d+1})$ such that $\{v_{i}:i\in I_{1}\}=\{1,2,\ldots,|I_{1}|\}$. Then $h\cdot F_{I}\cap F_{I}=F_{h\cdot I}\cap F_{I}=\emptyset$ for all $h\neq\operatorname{id}$ in $H$, and it follows that the action of $H$ on $K=\partial P$ is translative. Let $Q$ be the interior $H$-fixed lattice point $(\frac{p+1}{2},\ldots,\frac{p+1}{2})$. The claim now follows from Corollary 3.17 and Corollary 3.18.
@@ 3.4. The two dimensional case
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Assume that $d=2$. Then the following conditions are equivalent:
@@
The implication (1) $\Rightarrow$ (2) is clear. The implication (2) $\Rightarrow$ (3) follows from Theorem 1.4. We now verify the implication (3) $\Rightarrow$ (4). Suppose that (4) doesn’t hold. We need to show that $h^{*}(P,\rho;t)$ is not a polynomial. Firstly, consider case 4(a) above. Then the $g$-fixed locus $P^{g}$ of $P$ corresponds to the point $\{(\frac{1}{2},\frac{1}{2})\}$, and hence $\operatorname{Ehr}(P,\rho;t)(g)=\frac{1}{1-t^{2}}$. Since $\det(I-\widetilde{M}_{\mathbb{C}}t)(g)=(1-t)(1+t^{2})$, we have that $h^{*}(P,\rho;t)(g)=\frac{1+t^{2}}{1+t}$ has a pole at t=1t=-1. Secondly, in case 4(b), [Sta11a, Lemma 7.3] implies that $h^{*}(P,\rho;t)$ is not a polynomial.
@@
The subgroups of $\operatorname{GL}(2,\mathbb{Z})$, up to conjugacy, are classified [Vos65]. They are precisely the subgroups of the dihedral group $D_{4}$ of symmetries of $1,1]^{2}$, and the subgroups of the dihedral group $D_{6}$ of symmetries of the hexagon with vertices $\{\pm e_{1},\pm(e_{1}+e_{2}),\pm e_{2}\}$. Consider an element $g\neq\operatorname{id}$ in $G$. By Remark 2.9, the degree $2$ characteristic polynomial of the action of $g$ on $M_{\mathbb{C}}$ is a product of cyclotomic polynomials. Hence $\operatorname{ord}(g)\in\{2,3,4,6\}$ and if $1$ is an eigenvalue of $g$ acting on $M_{\mathbb{C}}$ then $\operatorname{ord}(g)=2$ (this also follows from the classification of subgroups of $\operatorname{GL}(2,\mathbb{Z})$). If $1$ is not an eigenvalue for the action of $g$ on $M_{\mathbb{C}}$, then by Lemma 2.5 and Lemma 2.6 either $\operatorname{ord}(g)\in\{2,4\}$ and $c\in\frac{1}{2}M$ or $\operatorname{ord}(g)=3$ and $c\in\frac{1}{3}M$. If $\operatorname{ord}(g)=4$ and $c\in\frac{1}{2}M$, then we see from the classification that we are in case 4(a), a contradiction. In particular, we have shown that $\operatorname{ord}(g)\in\{2,3\}$.
@@
Let $J$ be the set of lattice points in $P$ that are not vertices of $\mathcal{S}$. Since $\mathcal{S}$ equals the pulling refinement of $\mathcal{S}$ by $J$, we conclude that $J$ is the emptyset. Hence each facet $F$ of $\mathcal{S}$ is an empty polytope, in the sense that $F\cap M$ is precisely the set of vertices of $F$. On the other hand, it is well-known and straightforward to check that the only empty $2$-dimensional lattice polytopes in $\mathbb{R}^{2}$, up to unimodular equivalence, are the standard simplex and $0,1]^{2}$. Fix a facet $F$ of $\mathcal{S}$ that is not a simplex. Then we may assume that $M=\mathbb{Z}^{2}$ and $F=[0,1]^{2}$. Label the vertices of $F$: $u=0$, $v=e_{1}$, $w=e_{1}+e_{2}$, $x=e_{2}$.
@@ 4.1. Commutative algebra
First recall that any surjection of graded $\mathbb{C}G$-modules splits. Indeed, if $f:V\to W$ is a surjection of graded $\mathbb{C}G$-modules and $h^{\prime}:W\to V$ is a splitting of $f$ as a map of graded $\mathbb{C}$-vector spaces, then one may define $h:W\to V$ by $h(w)=\frac{1}{|G|}\sum_{g\in G}gh^{\prime}(g^{-1}w)$ for all $w$ in $W$.
Consider the graded subring $S=\mathbb{C}[x_{1},\ldots,x_{n}]$ of $R$ and let $\mathfrak{m}=(x_{1},\ldots,x_{n})$. Note that each $x_{i}$ has degree $r$. By Theorem 4.4, $S$ is isomorphic to a polynomial ring in $n$ variables. Since $R$ is Cohen-Macaulay, $R$ is a free $S$-module by Theorem 4.6.
Consider the surjection of graded $\mathbb{C}G$-modules $p:R\to R/J=R/\mathfrak{m}R$, and choose a splitting of graded $\mathbb{C}G$-modules $\sigma:R/J\to R$. Let $y_{1},\ldots,y_{s}$ be a $\mathbb{C}$-basis of $R/J$. Then $\sigma(y_{1}),\ldots,\sigma(y_{s})$ generate $R$ as an $S$-module by the graded version of Nakayama’s lemma (see [Eis95, Exercise 4.6a] and [AM69, Proposition 2.8]). Here $s$ is the rank of $R$ as a free $S$-module and hence $\sigma(y_{1}),\ldots,\sigma(y_{s})$ is a a basis of $R$ as an $S$-module [AM69, Exercise 3.15].
@@ 4.2. The principal $A$-determinant and nondegenerate polynomials
[GKZ94, Proposition 8.2.1] Let $A\subset M$ be a finite set with convex hull $P\subset M_{\mathbb{R}}$. With the notation above, consider the polynomial ring $\mathbb{Z}[z_{i,a}]$ for some formal variables $\{z_{i,a}:1\leq i\leq d+1,a\in A\}$. Then there is an irreducible polynomial (unique up to sign) $R_{A}(z_{i,a})\in\mathbb{Z}[z_{i,a}]$ called the $A$-resultant satisfying the following property:
Let $F_{1},\ldots,F_{d+1}$ be elements of $S_{P}$ with support contained in $A\times\{1\}$. For $1\leq i\leq d+1$, write $F_{i}=\sum_{a\in A}\lambda_{i,a}x^{(a,1)}$ for some $\lambda_{i,a}\in\mathbb{C}$. Then $F_{1},\ldots,F_{d+1}$ is a l.s.o.p. for $S_{P}$ if and only if $R_{A}(\lambda_{i,a})\neq 0$.
@@
In Definition 4.9, the statement given is different but equivalent to that in [GKZ94, Proposition 8.2.1]. Namely, [GKZ94] associate to $A$ the semigroup algebra $S_{A}:=\mathbb{C}[C_{P}\cap\widetilde{M}_{A}]$ and the (not necessarily normal) toric variety $\operatorname{Proj}S_{A}$. With the notation of Definition 4.9, the $A$-resultant in [GKZ94] is determined by the condition that $R_{A}(\lambda_{i,a})\neq 0$ if and only if the subvariety $\operatorname{Proj}(S_{A}/(F_{1},\ldots,F_{d+1}))$ of $\operatorname{Proj}S_{A}$ is empty.
We explain why this is equivalent to Definition 4.9 cf. the proof of [Bat93, Theorem 4.8] (stated in Theorem 4.18 below). First observe that $S_{A}\subset S_{P}$ and $S_{P}$ is integral over $S_{A}$. This follows, for example, by choosing a regular triangulation of $P$ with vertices in $A$, and then noting that every element in $C_{P}\cap\widetilde{M}$ is a sum of elements of $A\times\{1\}$ with nonnegative rational coefficients. In particular, $S_{A}$ has Krull dimension $d+1$ and $S$ is a finitely generated $S_{A}$-module by Lemma 4.1.
Now $\operatorname{Proj}(S_{A}/(F_{1},\ldots,F_{{d+1}}))$ is empty if and only if $\operatorname{Spec}(S_{A}/(F_{1},\ldots,F_{{d+1}}))$ is 0-dimensional. The latter holds if and only if $S_{A}/(F_{1},\ldots,F_{{d+1}})$ has Krull dimension 0 i.e. $F_{1},\ldots,F_{{d+1}}$ is a l.s.o.p. for $S_{A}$. By Corollary 4.5, $F_{1},\ldots,F_{{d+1}}$ is a l.s.o.p. for $S_{A}$ if and only if $F_{1},\ldots,F_{{d+1}}$ is a l.s.o.p. for $S_{P}$.
@@
[…]
@@
[GKZ94, 10.1.A] Let $A\subset M$ be a finite set with convex hull $P\subset M_{\mathbb{R}}$. With the notation above, consider the polynomial ring $\mathbb{Z}[z_{a}]$ for some formal variables $\{z_{a}:a\in A\}$. Fix a choice of basis $v_{1},\ldots,v_{d+1}$ of $\widetilde{N}_{A}$. For $a$ in $A$ and $1\leq i\leq d+1$, let $w_{i,a}:=\langle(a,1),v_{i}\rangle z_{a}$. Then $E_{A}(z_{a}):=R_{A}(w_{i,a})\in\mathbb{Z}[z_{a}]$.
@@
Definition 4.12 could alternatively have been described as follows. Fix a choice of basis $v_{1},\ldots,v_{d+1}$ of $\widetilde{N}_{A}$. The principal $A$-determinant $E_{A}(z_{a})\in\mathbb{Z}[z_{a}]$ is the (unique up to sign) polynomial satisfying the following property:
Let $F$ be an element of $S_{P}$ with support contained in $A\times\{1\}$, and write $F=\sum_{a\in A}\lambda_{a}x^{(a,1)}$ for some $\lambda_{a}\in\mathbb{C}$. For all $1\leq i\leq d+1$, write $\partial_{v_{i}}F=\sum_{a\in A}\lambda_{i,a}x^{(a,1)}$ for some $\lambda_{i,a}\in\mathbb{C}$. Then $E_{A}(\lambda_{a})=R_{A}(\lambda_{a,i})$.
In particular, with this notation, $\partial_{v_{1}}F,\ldots,\partial_{v_{d+1}}F$ is a l.s.o.p. for $S_{P}$ if and only if $E_{A}(\lambda_{a})\neq 0$.
@@
where $\operatorname{Vol}(P)$ is the normalized volume of $P$ with respect to $M$ and $\widetilde{M}:\widetilde{M}_{A}]$ is the index of the sublattice $\widetilde{M}_{A}$ of $\widetilde{M}$. Importantly, $E_{A,M}(\lambda_{a})=0$ if and only if $E_{A}(\lambda_{a})=0$.
@@
See [Stu94, (28)] for an analogous result to Theorem 4.19 for the AA-resultant.